### v1

@@ 3 Priority-Neutral Lattices Are Not Distributive
Recall that, like the classical set of stable matchings, the set of priority-neutral matchings has the mathematical structure of a lattice [Ren22]. In this section, we show that unfortunately, the $\mathsf{PN}$ lattice actually lacks much of the structure the lattice of stable matchings has. We construct a set of preferences and priorities such that $\mathsf{PN}$ is not be distributive, and that greatest lower bounds in $\mathsf{PN}$ are not given by coordinatewise minimums.
Figure 1 shows our construction. Figure 1 gives the priorities and preferences of the construction. We will show that Figure 1 displays the entire lattice of priority-neutral matchings for this instance, where $\mu\leq\nu$ if any only if there is a path from $\mu$ to $\nu$ consisting of edges ordered in their upward direction. The matchings in Figure 1 are described in terms of the rotations given in Figure 1, as discussed in Section 2.1. For brevity, we omit the “base” matching $\mu_{0}$ from our notation for different matchings; each time we write $\alpha_{1}\ldots\alpha_{k}$ through this section for rotations $\alpha_{1},\ldots,\alpha_{k}$, we mean the matching $\alpha_{1}\ldots\alpha_{k}\mu_{0}$.888 For example, the matching denoted $\alpha$ equals $[(h_{A},d_{B}),(h_{B},d_{A}),(h_{C},d_{C}),(h_{D},d_{D}),(h_{E},d_{E}),(h_{F},d_{F}),(h_{Z},d_{Z})]$, and the matching $\beta\gamma$ equals $[(h_{A},d_{A}),(h_{B},d_{B}),(h_{C},d_{D}),(h_{D},d_{C}),(h_{E},d_{F}),(h_{F},d_{E}),(h_{Z},d_{Z})]$.
@@
We start by giving an high-level description of the important properties of the example in Figure 1. The crucial fact about this construction is that $\alpha\beta$ and $\beta\gamma$ are priority-neutral, but $\beta$ is not priority-neutral. This will immediately imply that $\mathsf{PN}$ is not closed under coordinatewise minimums; indeed, $\alpha\beta$ and $\beta\gamma$ serve a role analogous to $\mu_{L}$ and $\mu_{R}$ in 2.6. We found a construction with this property by combining two things. First, there is an applicant $d_{Z}$ can be Pareto-improved in $\mu_{0}$, the institution-optimal stable matching; however, either of the rotations $\alpha$ or $\gamma$ make this Pareto-improvement no longer possible.999 The most involved part of the proof below involves arguing that in $\alpha\beta$ and $\beta\gamma$, not only is $d_{Z}$ not Pareto-improvable, but $d_{Z}$ has no priority-correcting adjustment (2.4). Second, $d_{Z}$’s priority is violated in $\beta$, and additionally $d_{Z}$ can Pareto-improve in $\beta$ (as in $\mu_{0}$), showing that $\beta\notin\mathsf{PN}$. As we argue below, all of the other relevant combinations of rotations will be priority-neutral matchings, showing that $\mathsf{PN}$ is given by the lattice in Figure 1.
@@
To begin, we check that the stable matching lattice, which is always contained in $\mathsf{PN}$ by Theorem 2.5, is given by the green-shaded matchings in Figure 1. First, the institution-optimal matching equals $\mu_{0}$, as defined in Figure 1, since all institutions have distinct top-priority applicants equal to their match in $\mu_{0}$. Second, one can run applicant-proposing deferred acceptance to see that $\alpha\gamma$ is the applicant-optimal stable matching. Third, one can verify directly that the matchings $\alpha$ and $\gamma$ have no priority violations, and are thus stable; since no other distinct matchings $\mu^{\prime}$ exist with $\mu_{0}<\mu^{\prime}<\alpha\gamma$, these green shaded matchings are all stable matchings.101010 In full detail, in $\alpha$, all institutions except for $h_{A}$ and $h_{B}$ are matched to their top-priority applicants; while $d_{E}$ prefers $h_{A}$ to their match in $\alpha$, we have $d_{B}\succ_{h_{A}}d_{E}$, so $\alpha$ is stable. Similarly, in $\gamma$, all institutions except for $h_{E}$ and $h_{F}$ are matched to their top-priority applicants; while $d_{Z}$ prefers $h_{E}$ to their match in $\gamma$, we have $d_{F}\succ_{h_{E}}d_{Z}$, and $\gamma$ is stable. Alternatively, the methods of [GI89] can be used to directly compute the set of rotation representing the stable matching lattice (which are $\alpha,\gamma$ with no predecessor relations).
@@
Now, three relevant matchings remain which we have not yet considered: $\beta$, $\alpha\beta$, and $\beta\gamma$. Since these are the only not-yet-considered matchings $\mu$ such that $\mu_{0}<\mu<\alpha\beta\gamma$, we know that all non-stable matchings in $\mathsf{PN}$ must be in $\{\beta,\alpha\beta,\beta\gamma\}$ by Theorem 2.5. We will proceed by showing that $\beta\notin\mathsf{PN}$, but $\alpha\beta\in\mathsf{PN}$ and $\beta\gamma\in\mathsf{PN}$.
@@
$\alpha\beta:$ $h_{A}$ $h_{B}$ $h_{C}$ $h_{D}$ $h_{E}$ $h_{F}$ $h_{Z}$ $d_{B}$ $d_{A}$ $d_{D}$ $d_{C}$ $d_{E}$ $d_{F}$ $d_{Z}$     $\beta\gamma:$ $h_{A}$ $h_{B}$ $h_{C}$ $h_{D}$ $h_{E}$ $h_{F}$ $h_{Z}$ $d_{A}$ $d_{B}$ $d_{D}$ $d_{C}$ $d_{F}$ $d_{E}$ $d_{Z}$
@@
We can now use this example to give the main result of this paper. This resolves an open question of [Ren22]: the greatest lower bound of two matchings in $\mathsf{PN}$ need not be their coordinatewise minimum. It also shows that, crucially, the $\mathsf{PN}$ lattice is not distributive, and hence cannot be represented by a partial order on a set of rotations, as for the stable lattice.
Theorem 3.2 ($\mathsf{PN}$ need not be distributive).
$\mathsf{PN}$ may not be a distributive lattice. Moreover, the greatest lower bound of two matchings in $\mathsf{PN}$ may not be their coordinatewise minimum.
@@
where we recall that $\wedge_{\mathsf{PN}}$ is the greatest lower bound of two matchings within $\mathsf{PN}$, and $\wedge_{\mathsf{L}}$ is their coordinatewise minimum. To prove the first sentence, we build on the above fact (used in equality (*)) and observe that
@@
4 Not All Lattices Arise From Priority-Neutral Matchings
In this section, we show that not all lattices arise as the set of priority-neutral matchings for some priorities and preferences. To prove this result, we give a novel structural property of the $\mathsf{PN}$ lattice, how it relates to the legal lattice $\mathsf{L}$ of [EM20], and about certain special types of chains of matchings, i.e. sequences $\mu_{1}\leq\mu_{2}\leq\ldots\leq\mu_{k}$. Then, we use these properties to rule out a specific lattice; as we discuss below, this lattice is in fact a classic example known as $M_{3}$ in the order theory literature.
4.1 Structural properties

### v2

@@ 3 Priority-Neutral Lattices Are Not Distributive
In this section, we show that the $\mathsf{PN}$ lattice lacks much of the tractable structure of the stable lattice. We give a self-contained proof that $\mathsf{PN}$ is not distributive, and that greatest lower bounds in $\mathsf{PN}$ are not given by coordinatewise minimums.
We prove that the above results via an explicit example, presented in Figure 2. Figure 2 gives the priorities and preferences of the construction. As we show below, Figure 2 displays the $\mathsf{PN}$ lattice for this instance, where $\mu\leq\nu$ if any only if $\mu$ is below $\nu$ in a path in the diagram. Figure 2 displays $\mathsf{IPDA}$ and the set of rotations used to describe $\mathsf{PN}$, as discussed in Section 2.1. For brevity, through this section we write $\alpha_{1}\ldots\alpha_{k}$ to mean $\alpha_{1}\ldots\alpha_{k}\mathsf{IPDA}$. [^8] For example, the matching denoted $\alpha$ equals $\{(h_{A},d_{B}),(h_{B},d_{A}),(h_{C},d_{C}),(h_{D},d_{D}),(h_{E},d_{E}),(h_{F},d_{F}),(h_{Z},d_{Z})\}$, and the matching $\beta\gamma$ equals $\{(h_{A},d_{A}),(h_{B},d_{B}),(h_{C},d_{D}),(h_{D},d_{C}),(h_{E},d_{F}),(h_{F},d_{E}),(h_{Z},d_{Z})\}$.
@@
Note: The vertical dots indicate that the remainder of the preference/priority list is irrelevant.
@@
The crucial fact about this construction is that $\alpha\beta$ and $\beta\gamma$ are priority-neutral, but $\beta$ is not priority-neutral. This will immediately imply that $\mathsf{PN}$ is not closed under coordinatewise minimums (since indeed $\beta$ is the coordinatewise minimum of $\alpha\beta$ and $\beta\gamma$). Two important things are used to guarantee this property. First, there is an applicant $d_{Z}$ can be Pareto-improved in $\mathsf{IPDA}$; however, either of the rotations $\alpha$ or $\gamma$ make this Pareto-improvement no longer possible. [^9] The most involved part of the proof below involves arguing that in $\alpha\beta$ and $\beta\gamma$, not only is $d_{Z}$ not Pareto-improvable, but $d_{Z}$ has no priority-correcting adjustment (2.4). Second, $d_{Z}$’s priority is violated in $\beta$, and additionally $d_{Z}$ can Pareto-improve in $\beta$ (as well as $\mathsf{IPDA}$), showing that $\beta\notin\mathsf{PN}$. As we argue below, all of the other relevant combinations of rotations will be priority-neutral matchings, showing that $\mathsf{PN}$ is given by the lattice in Figure 2.
@@
To begin, we check that the stable matching lattice, which is always contained in $\mathsf{PN}$ by Theorem 2.5, is given by the green-shaded matchings in Figure 2. First, the institution-optimal matching equals $\mu_{0}$, as defined in Figure 2, since all institutions have distinct top-priority applicants equal to their match in $\mu_{0}$. Second, one can run applicant-proposing deferred acceptance to see that $\alpha\gamma$ is the applicant-optimal stable matching. Third, one can verify directly that the matchings $\alpha$ and $\gamma$ have no priority violations, and are thus stable; since no other distinct matchings $\mu^{\prime}$ exist with $\mu_{0}<\mu^{\prime}<\alpha\gamma$, these green shaded matchings are all stable matchings. [^10] In full detail, in $\alpha$, all institutions except for $h_{A}$ and $h_{B}$ are matched to their top-priority applicants; while $d_{E}$ prefers $h_{A}$ to their match in $\alpha$, we have $d_{B}\succ_{h_{A}}d_{E}$, so $\alpha$ is stable. Similarly, in $\gamma$, all institutions except for $h_{E}$ and $h_{F}$ are matched to their top-priority applicants; while $d_{Z}$ prefers $h_{E}$ to their match in $\gamma$, we have $d_{F}\succ_{h_{E}}d_{Z}$, and $\gamma$ is stable. Alternatively, the methods of [GI89] can be used to directly compute the set of rotation representing the stable matching lattice (which are $\alpha,\gamma$ with no predecessor relations).
@@
Now, three relevant matchings remain which we have not yet considered: $\beta$, $\alpha\beta$, and $\beta\gamma$. Since these are the only not-yet-considered matchings $\mu$ such that $\mu_{0}<\mu<\alpha\beta\gamma$, we know by Theorem 2.5 that all non-stable matchings in $\mathsf{PN}$ must be in $\{\beta,\alpha\beta,\beta\gamma\}$. We will proceed by showing that $\beta\notin\mathsf{PN}$, but $\alpha\beta\in\mathsf{PN}$ and $\beta\gamma\in\mathsf{PN}$.
@@
$\alpha\beta:$ $h_{A}$ $h_{B}$ $h_{C}$ $h_{D}$ $h_{E}$ $h_{F}$ $h_{Z}$ $d_{B}$ $d_{A}$ $d_{D}$ $d_{C}$ $d_{E}$ $d_{F}$ $d_{Z}$     $\beta\gamma:$ $h_{A}$ $h_{B}$ $h_{C}$ $h_{D}$ $h_{E}$ $h_{F}$ $h_{Z}$ $d_{A}$ $d_{B}$ $d_{D}$ $d_{C}$ $d_{F}$ $d_{E}$ $d_{Z}$ .
@@
We can now prove our first theorem. The above example shows that the greatest lower bound of two matchings in $\mathsf{PN}$ need not be their coordinatewise minimum, resolving an open question of [Ren22]. It also shows that the $\mathsf{PN}$ lattice is not distributive.
Theorem 3.2 ($\mathsf{PN}$ Is Not Distributive).
$\mathsf{PN}$ is not a distributive lattice. Moreover, the greatest lower bound of two matchings in $\mathsf{PN}$ is not their coordinatewise minimum.
@@
where we write $\wedge_{\mathsf{PN}}$ to denote the greatest lower bound of two matchings within $\mathsf{PN}$, and $\wedge_{\mathsf{L}}$ to denote their coordinatewise minimum. To prove the first sentence, we build on the above fact (used in equality (*)) and observe that
@@
4 Building Up To Our Main Result
In this section, we build up to our main result. First, in Section 4.1, we recall more-detailed preliminaries and prior results which will be important for our analysis. Second, in Section 4.2, we prove new structural results of $\mathsf{PN}$, particularly a novel “good chain” property. Third, in Section 4.3, we prove a partial characterization, namely a necessary condition that all $\mathsf{PN}$ lattices must satisfy, as a tool to build towards our main characterization and illustrate the intricacies involved.
4.1 Additional Preliminaries: Rotation DAG and Legal Lattice
4.1.1 Rotation DAG Representation of $\mathsf{Stab}$
Following the classical theory of stable matchings, we represent matching lattices via subsets of rotations (introduced as notation in Section 2.1). We recall here all the properties and definitions we need, but defer full coverage of some topics (such as algorithms for actually finding the rotation representation) to the standard reference [GI89]. See also [CT19] or [EIV23, Section 1.4]. [^11] While the notion in our definitions are identical to classical representations, we phrase the definitions with slightly more data (such as having a fixed ordering on the rotations that gives a topological sort of the DAG). We find these extra pieces of data conceptually helpful when we generalize this construction in later sections, since they help express additional conditions on these generalized constructions.
Recall that a rotation $\rho=[(h_{1},d_{1}),\allowbreak\ldots,\allowbreak(h_{k},d_{k})]$, where each $h_{i}\in\mathcal{I}$ and $d_{i}\in\mathcal{A}$, is used to represent the difference between two matchings. If $\mu$ has $\mu(h_{i})=d_{i}$ for each $i$, we say $\rho$ is valid at $\mu$. In this case, we let $\alpha\mu$, called the elimination of $\alpha$ from $\mu$, denote the matching which differs from $\mu$ only by matching $d_{i}$ to $h_{i+1}$ for $i=1,\ldots,k$ modulo $k$. (We also denote rotations with letters such as $\alpha,\beta,\gamma$ in specific examples.)
Definition 4.1 (Rotation DAG of distributive lattices).
For some set of applicants $\mathcal{A}$ and institutions $\mathcal{I}$, a rotation DAG is a tuple $G=(\mu,R,P)$, defined as follows. First, $\mu$ is some matching, referred to as the “base matching.” Second, $R$ is some set of rotations $R=\{\rho_{1},\ldots,\rho_{N}\}$, equipped with a fixed ordering $\rho_{1},\ldots,\rho_{N}]$. Third, $P$ is a set of predecessor relationships over $R$. Each such predecessor relationship is a pair $p=(\rho_{i},\rho_{j})\in R\times R$, where $i>j$; we also think of such a $p$ as an edge $\rho_{i}\to\rho_{j}$, and in this case we say that $\rho_{j}$ is a predecessor of $\rho_{i}$ (and the fact that we always have $i>j$ means that $(R,P)$ forms a DAG).
Consider some subset $X\subseteq R$ and some predecessor relationship $p=(r_{0},r_{1})\in P$. We say that $p$ is violated by $X$ if $r_{0}\in X$, but $r_{1}\notin X$. [^12] This terminology is inspired by the fact that $S\subseteq R$ represents a matching $\mu$, and whenever some predecessor relationship $p$ is violated by $S$, this means that some applicant’s priority is violated in $\mu$.
Define the (abstract) distributive lattice, denoted $D_{\mathsf{abs}}(G)$, to be the collection of subsets of $R$, ordered by set inclusion, defined as follows:
$D_{\mathsf{abs}}(G)=\left\{X\subseteq R\ \big|\ \text{No $p\in P$ is violated by $X$}\right\}.$
Define the (matching) distributive lattice, denoted $D(G)$, as follows. For every $X\in D_{\mathsf{abs}}(G)$, we require that if $X=\{\rho_{i(1)},\ldots\rho_{i(k)}\}$, where $i(1)<\ldots<i(k)$, then $\rho_{i(j+1)}$ is valid at $\rho_{i(j)}\ldots\rho_{i(1)}\mu$ for each $j=0,\ldots,k-1$. Then, define $\mathsf{match}(X)=\rho_{i(k)}\ldots\rho_{i(1)}\mu$. We set
$D(G)=\left\{\mathsf{match}(X)\ \big|\ \text{$X\subseteq R$, and no $p\in P$ is violated by $X$}\right\}.\qed$
Note that any set $R$ and predecessor DAG on $R$ can define an abstract lattice $D_{\mathsf{abs}}(G)$ on subsets of $R$. The matching lattice $D(G)$ then specifies, for each $X\in D_{\mathsf{abs}}(G)$, a corresponding matching $\mathsf{match}(X)$ which arises by eliminating all of the rotations in $X$ from the “base” matching $\mu$. For a matching $\mu\in D(G)$, we also define $\mathsf{rots}(\mu)=X\in D_{\mathsf{abs}}(G)$ to be the unique $X\subseteq R$ such that $\mathsf{match}(X)=\mu$.
The classical representation of $\mathsf{Stab}$ is as follows:
Theorem 4.2 ([GI89]).
There exists a set of rotations $R$ and predecessors $P$, such that if $G=(\mathsf{IPDA},R,P)$, then $D(G)=\mathsf{Stab}$.
Additionally, [GI89, CT19] show that we can identify the rotations and predecessors as follows. Recall that in any lattice, we say $\mu$ covers $\nu$, denoted $\mu\gtrdot\nu$, if $\mu>\nu$ and for all $\xi$ with $\mu\geq\xi\geq\nu$, we have $\xi=\mu$ or $\xi=\nu$. The set of rotations $R$ is the set of all $\rho$ such that there exists a $\mu\in\mathsf{Stab}$ where $\rho$ is valid in $\mu$, and $\rho\mu\in\mathsf{Stab}$, and where $\rho\mu$ covers $\mu$ in $\mathsf{Stab}$; intuitively, this represents the “minimal possible differences” between two stable matchings. (Note in particular that rotations always improve that match of applicants, and worsen the match of institutions.) In this case, we say that $\rho$ is exposed in $\mu$. [^13] Note that even if $\rho$ is valid in $\mu$, it may not be the case that $\rho$ is exposed in $\mu$. The predecessors $P$, which come in two different types, can be identified as follows:
  1. 1.
    Rotation $\rho_{1}$ is a type 1 predecessor of rotation $\rho_{2}$ if there exists a $(d,h)\in\mathcal{A}\times\mathcal{I}$ such that $\rho_{1}$ moves $d$ to $h$, and $\rho_{2}$ moves $d$ above $h$. In other words, we have:
    $\rho_{2}=[\ldots,(h,d),\ldots]$ for all such rotations.
     $\downarrow\text{ (type 1) }$
    $\rho_{1}=[\ldots,(\cdot,d),(h,\cdot),\ldots]$
  2. 2.
    Rotation $\rho_{1}$ is a type 2 predecessor of rotation $\rho_{2}$ if there exists a $(d,h)\in\mathcal{A}\times\mathcal{I}$ such that $\rho_{1}$ moves $d$ from below $h$ to above $h$, and $\rho_{2}$ moves $h$ from above $d$ to below $d$. In other words, we have:
    $\rho_{2}=[\ldots,(\cdot,d_{\text{down}}),(h,d_{\text{up}}),\ldots]$ if we have $h_{\text{up}}\succ_{d}h\succ_{d}h_{\text{down}}$,
     $\downarrow\text{ (type 2) }$ and also $d_{\text{up}}\succ_{h}d\succ_{h}d_{\text{down}}$.
    $\rho_{1}=[\ldots,(h_{\text{down}},d),(h_{\text{up}},\cdot),\ldots]$
Intuitively, type 1 predecessor relations are needed because applicants can only “switch stable partners” in the units specified by rotations, and because if such a $d$ is improved above $h$ by $\rho_{2}$, then $d$ must have already be improved to $h$ by $\rho_{1}$. It is easy to see why type 2 predecessor relations are needed: if we eliminate such a $\rho_{2}$ without eliminating $\rho_{1}$, then $d$’s priority at $h$ is violated, so the matching is not stable. For example, the stable matching lattice in Figure 1 is represented by rotations $\alpha=[(h_{A},d_{A}),(h_{B},d_{B})]$ and $\beta=[(h_{C},d_{C}),(h_{D},d_{D})]$, where $\alpha$ is a type 2 predecessor of $\beta$. See [GI89] or [CT19] for details. We build on the above concrete representation of the rotations and predecessors in Section 5 when we prove our full characterization.
4.1.2 The Legal Matching Lattice
While our main focus is on the priority-neutral lattice $\mathsf{PN}$, we make use of another interesting generalization of $\mathsf{Stab}$ known as the legal lattice [EM20] (already introduced in Footnote 3), which will aid us greatly in our analysis. A set $\mathsf{L}$ of matchings is called legal if
[…]
[EM20] prove that there exists a unique legal set $\mathsf{L}$, which always contains $\mathsf{Stab}$, and that the applicant-optimal element of $\mathsf{L}$ is $\mathsf{EADAM}$. Thus, $\mathsf{PN}$ and $\mathsf{L}$ have the same topmost element; moreover, [Ren22] proves that $\mathsf{PN}\subseteq\mathsf{L}$, so these two lattices are closely related.
In contrast to $\mathsf{PN}$, however, $\mathsf{L}$ has the same tractable structure as $\mathsf{Stab}$. [FZ22] formalizes this by showing that every legal lattice can be represented as the stable lattice on some altered set of preferences and priorities. This allows us to greatly simply our analysis (intuitively, by representing $\mathsf{PN}$ as a subset of $\mathsf{L}$, where we use the “off the shelf” rotation DAG of $\mathsf{L}$, and specify which legal matchings are not priority-neutral). In summary, the main known results are the following:
Theorem 4.3 ([EM20, Ren22, FZ22]).
For convenience, in most of our analysis we actually use the sub-lattice $\mathsf{K}$ of $\mathsf{L}$ which also has the same bottommost element as $\mathsf{PN}$, namely $\mathsf{IPDA}$. [^14] To see that $\mathsf{L}$ does not in general have the same lowest elements as $\mathsf{PN}$, observe that the definition of $\mathsf{L}$ is symmetric in the roles of applicants and institutions, and thus $\mathsf{L}$ always contains the institution-side analogue of $\mathsf{EADAM}$ (which is a different matching from $\mathsf{IPDA}$). In other words, we define $\mathsf{K}=\{\mu\in\mathsf{L}\ \mid\ \mu\geq\mathsf{IPDA}\}$; note that $\mathsf{PN}\subseteq\mathsf{K}$. It is easy to see that $\mathsf{K}$ is also distributive. See Figure 3 for an illustration of all matching lattices we consider.
Figure 3: Illustration of matching lattices.
Notes: As the figure shows, $\mathsf{Stab}\subseteq\mathsf{PN}\subseteq\mathsf{K}\subseteq\mathsf{L}$. Each of $\mathsf{Stab},\mathsf{K}$, and $\mathsf{L}$ are always distributive lattices, and can be represented with a predecessor DAG on a set of rotations, and hence we illustrate each of these with a straight-bordered shape. In contrast, as shown in Section 3, $\mathsf{PN}$ is not distributive. The applicant optimal / pessimal matchings in each lattice are labeled within parentheses, where the matching $\underline{\mu}$ denotes the result of running $\mathsf{EADAM}$ on the instance produced by interchanging the roles of the applicants and the institutions.
The fact that the lattice $\mathsf{K}$ has a rotation DAG representation is guaranteed by the following:
Lemma 4.4.
There always exists a rotation DAG, call it $G$, such that $D(G)=\mathsf{K}$.
Proof.
Fix some instance with preferences and priorities $Q$. The main result of [FZ22] is that there exists an altered instance $\widetilde{Q}$ such that $\mathsf{L}$ equals the set of stable matchings under $\widetilde{Q}$. Thus, lattice $\mathsf{L}$ under $Q$ has a predecessor DAG representation $(\underline{\mu},R^{\prime},P^{\prime})$, where $\underline{\mu}$ the the result of running $\mathsf{EADAM}$ while interchanging the role of applicants and institutions, which is equal to the representation of the stable lattice under $\widetilde{Q}$.
To transform this rotation DAG into an $(\mathsf{IPDA},R,P)$ which represents $\mathsf{K}$, it suffices to simply define $R$ by removing from $R^{\prime}$ the set of rotations corresponding to $\mathsf{IPDA}$ (and setting $P$ to remove all elements of $P^{\prime}$ containing removed elements of $R^{\prime}$). After doing this operation, it is straightforward to verify that $D(\mathsf{IPDA},R,P)$ contains exactly those elements of $D(\underline{\mu},R^{\prime},P^{\prime})$ that weakly dominate $\mathsf{IPDA}$, as desired. ∎
4.2 Structural Results for $\mathsf{PN}$ and the Good Chain Property
We now begin to reason about the structure of $\mathsf{PN}\subseteq\mathsf{K}$. The main structural result we establish here is that, while many matchings in $\mathsf{K}$ may not be in $\mathsf{PN}$, there always exists a chain $\mathsf{IPDA}=\mu_{1}<\mu_{2}<\ldots<\mu_{N}=\mathsf{EADAM}$ in $\mathsf{PN}$, such that each $\mu_{i+1}$ differs from $\mu_{i}$ by a rotation in $\mathsf{K}$. To prove this, we adapt and prove additional properties of several notions of matching chains from prior work, particularly [TY14, FZ22, Ren22].